OpenStudy (anonymous):

Solve for x . √4x+5=x

OpenStudy (campbell_st):

subtract 5 from both sides \[\sqrt{4x} = x - 5\] next square both sides to remove the radical \[4x = (x -5)^2\] expand and simplify \[4x = x^2 - 10x + 25\] subtract 4x from both sides \[x^2 - 14x + 25 = 0\] use the general quadratic formula with a = 1, b = -14 and c = 25 \[x = \frac{14 \pm \sqrt{196 - 4 \times1\times25}}{2\times1}\] \[x = \frac{14 \pm 4\sqrt{6}}{2} ... x = 7\pm2\sqrt{6}\]